OpenStudy (anonymous):

What is the simplified form of x-3/x^2+x-12 • x+4/x^2+8x+16 ?

OpenStudy (anonymous):

\[\frac{ x-3 }{ x ^{2}+x-12}\times \frac{ x+4 }{ x ^{2}+8x+16 }\]Is this the correct question?

OpenStudy (anonymous):

what are you trying to say with the dot?

OpenStudy (anonymous):

its times with the X. ands thats the correct question

OpenStudy (anonymous):

The first thing I see is you can factor the denominators and if done correctly the numerators will cancel. Try this first.

OpenStudy (anonymous):

i dont know how to do that, as in like, i need visual